OpenStudy (anonymous):

Verify the following identity : cotx – tanx x = [(4 cos ^(2)x – 2) / (sin 2x) ]

OpenStudy (anonymous):

@Meepi can we discuss the question???

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

oh okay.....

OpenStudy (anonymous):

i started out like this, but i am stuck! starting with the LEFT HAND SIDE, \[\frac{ 4\cos^2 x - 2 }{ \sin 2x }\] \[\frac{ 2 (2\cos ^2 x - 1) }{ 2\cos x \sin x }\] \[\frac{ \cos^2 x - 1 }{ \cos x \sin x }\]

OpenStudy (anonymous):

I have the same thing lol

OpenStudy (anonymous):

i made a mistake.....for the 3rd step, it is supposed to be \[\frac{ 2\cos^2 x- 1 }{ \cos x \sin x }\]

OpenStudy (anonymous):

what about if we simplify (2cos^2x - 1) to cos 2x?

OpenStudy (anonymous):

i thought of doing that but won't it make it more complicated??

OpenStudy (anonymous):

yeah your right

OpenStudy (anonymous):

i think that would work.....hold on...i am trying something

OpenStudy (anonymous):

alright thanks :)

OpenStudy (anonymous):

\[\frac{ \cos 2x }{ \cos x \sin x }\] \[\frac{ \cos^2x-\sin^2x }{ \cos x \sin x }\] \[\frac{ \cos^2x }{ \cos x \sin x } - \frac{ \sin^2 x }{ \cos x \sin x }\]

OpenStudy (anonymous):

do you get what i am trying to do??

OpenStudy (anonymous):

and can you continue from there??

OpenStudy (anonymous):

yeah I get what your trying to do

OpenStudy (anonymous):

I got it!!!

OpenStudy (anonymous):

yh.....good!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

do u need for me to show u what I did?

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

okay...sure you are welcome...

OpenStudy (anonymous):

cos ^ 2x / ( cos x sin x ) - (sin ^2x / (cosx sin x )) thats what your wrote but if u look carefully they cancel out.. u end up with ( cosx / sin x) - (sin x/ cos x) = cot x - tan x

OpenStudy (anonymous):

yh.....that is it!....(:

OpenStudy (anonymous):

awesome thank u!

OpenStudy (anonymous):

you are welcome!....